# Gravitational wave polarization modes and the kinematical tensors in general relativity and beyond

### v1

@@ Acknowledgments
The authors gratefully acknowledge valuable literature suggestions from E. Ayón-Beato and D. Flores-Alfonso, and worthwhile discussions with H. Quevedo. Some of the calculations done in this work were validated using Mathematica’s package xTensor [29]. C. M. acknowledges support from Secihti graduate fellowship. P. A. S. work was supported by Secihti postdoctoral fellowship, contract I1200/320/2022 Add. 2024.
@@ Appendix A Polarization modes in terms of the Newman-Penrose curvature scalars
To ease comparison of our results with others reported in the literature, we write here the polarization modes in terms of the Newman-Penrose curvature scalars [30]. In terms of the 4-velocity field $u$ of a cloud of test particles and the spatial propagation direction $n$, as seen in the instantaneous rest frame of the particles, of a non-necessarily weak gravitational wave [cf. (11)], one can introduce a complex null tetrad $(k,l,m,\overline{m})$, namely one such that $g_{ab}=-2k_{(a}l_{b)}+2m_{(a}\overline{m}_{b)}$, defined by
@@
Then, the polarization modes (13) can be written as
@@
${K^{\mathrm{v}}}_{a}=\frac{1}{\sqrt{2}}\real\left(\left[S_{bc}k^{b}\overline{m}^{c}-S_{bc}l^{b}\overline{m}^{c}+2\left(\overline{\Psi_{1}}-\Psi_{3}\right)\right]m_{a}\right),$ (57c)
${K^{\mathrm{tt}}}_{ab}=\real\left(\left[S_{cd}\overline{m}^{c}\overline{m}^{d}-\left(\overline{\Psi_{0}}+\Psi_{4}\right)\right]m_{a}m_{b}\right),$ (57d)

### v2

@@ Acknowledgments
The authors gratefully acknowledge valuable suggestions from E. Ayón-Beato, D. Flores-Alfonso, and G. Aldaz, worthwhile discussions with H. Quevedo, as well as insightful inquiries from an anonymous referee that improved this paper. Some of the calculations done in this work were validated using Mathematica’s package xTensor [40].
Funding
C. M. acknowledges support from Secihti graduate fellowship. P. A. S. work was supported by Secihti postdoctoral fellowship, contract I1200/320/2022 Add. 2024.
Data Availability Statement
Not applicable.
Code Availability Statement
Not applicable.
@@ Appendix A Polarization modes in terms of the Newman-Penrose curvature scalars
To ease comparison of our results with others reported in the literature, we write here the polarization modes in terms of the irreducible parts of the Riemann tensor. Using the irreducible decomposition of the Riemann tensor (cf., e.g., §3.5 of [41]) on (13), supplemented with (9), allows one to write
$K^{\mathrm{v}}_{a}$ $=\frac{1}{2}S_{bc}{p_{a}}^{b}n^{c}-C_{bcde}{p_{a}}^{b}u^{c}n^{d}u^{e},$ (54c)
$K^{\mathrm{tt}}_{ab}$ $=\frac{1}{2}S_{cd}{p_{a}}^{c}{p_{b}}^{d}+\frac{1}{4}\left(S_{cd}n^{c}n^{d}-S_{cd}u^{c}u^{d}\right)p_{ab}$
$\quad-C_{cdef}{p_{a}}^{c}u^{d}{p_{b}}^{e}u^{f}-\frac{1}{2}C_{cdef}u^{c}n^{d}u^{e}n^{f}p_{ab},$ (54d)
where $S_{ab}=R_{ab}-(R/4)g_{ab}$ is the traceless Ricci tensor; and, recalling, $u$ is the 4-velocity field of the cloud of test particles, and $n$ is the spatial propagation direction, as seen in the instantaneous rest frame of the particles, of a non-necessarily weak gravitational wave [cf. (11)]. It stands out that the scalar curvature $R$ does not contribute to the vector and transverse tensor modes.
One can obtain analogous expressions in terms of the Newman-Penrose curvature scalars [42]. To accomplish this, introduce a complex null tetrad $(k,l,m,\overline{m})$, namely one such that $g_{ab}=-2k_{(a}l_{b)}+2m_{(a}\overline{m}_{b)}$, defined by
@@
Then, the polarization modes can be written as
@@
$K^{\mathrm{v}}_{a}=\frac{1}{\sqrt{2}}\left(\left[S_{bc}k^{b}\overline{m}^{c}-S_{bc}l^{b}\overline{m}^{c}+2\left(\overline{\Psi_{1}}-\Psi_{3}\right)\right]m_{a}\right),$ (56c)
$K^{\mathrm{tt}}_{ab}=\left(\left[S_{cd}\overline{m}^{c}\overline{m}^{d}-\left(\overline{\Psi_{0}}+\Psi_{4}\right)\right]m_{a}m_{b}\right),$ (56d)